# Data-Juicer Sandbox: A Comprehensive Suite for Multimodal Data-Model Co-development

### v1

@@ A.2 Performance Metrics
where $N$ is the number of involved metrics, $s_{i}$ is the score of $i$-th model measurement metric, $s^{\prime}_{i}$ is the corresponding score gained by the baseline model. Below are the specific evaluation metrics involved in this study.
@@
A.3 Operator Descriptions
The study involves a total of 31 operators (OPs) from Data-Juicer (Chen et al., 2024a), with their corresponding statistics and detailed descriptions provided in Table 3.
@@ A.3 Operator Descriptions
image_text _similarity_filter text, image CLIP Image-Text Similarity Image-text feature cosine similarity based on a BLIP model
@@
Appendix B Additional Experimental Results
Figure 5: The VBench Leaderboard as of July 16, 2024, illustrating the rank 1 achievement empowered by our data-model co-development workflow.
B.1 Complete Operator Ranking
@@
In Table 4, we present complete numeric results conducted on individual OP experiments (Section 5.1), from which we can discern some more detailed observations.
In image-to-text generation, it is preferable for the input of training images to align as closely as possible with the original configuration of the vision tower, such as training dimensions (height, width, sizes). Additionally, CLIP similarity scores tends to be more reliable than BLIP similarity scores. The BLIP similarity does not show much distinction and paradoxically, a lower similarity often results in better performance, which defies common sense. Images with excessively high aesthetic quality may offer limited assistance in feature alignment, while watermarks might have certain impacts on the OCR performance of the model.
@@
B.2 Correlation Analysis
@@
In order to investigate the intrinsic relationships between OPs and to aid our recipe formulation, we observe relevance from following two perspectives.
@@
Figure 6: Pearson correlation coefficient of OP statistics from image-to-text generation.
@@
Figure 7: Pearson correlation coefficient of OP statistics from text-to-video generation.
@@
Firstly, we adopt the most direct approach by examining the Pearson correlation coefficient between the statistics of OPs, as shown in Figures 6 for the image-to-text generation and Figures 7 for text-to-video generation. Intuitively, the associations between the statistics of OPs utilized in the image-to-text generation appear to be much stronger compared to those in the text-to-video generation. For example, in image-to-text generation, phrase grounding recall shows a strong positive correlation with text perplexity and special character ratio, while exhibiting a strong negative correlation with alphanumeric ratio, language score, text action number, stopword ratio, and text length. Meanwhile, in text-to-video generation, we only observe relationships between some of the purely textual OPs, with video-related operators being mostly orthogonal to the others. Therefore, for the image-to-text generation, we divide OPs into three categories based on the correlation of their statistics and select the optimal OP from each category to create combinations. However, this method does not seem quite appropriate for text-to-video generation, due to the sparser correlations observed.
@@
Figure 8: Pearson Correlation Coefficient of each dimension in TextVQA, MMBench, and MME
Figure 9: Pearson Correlation Coefficient of Evaluation Metrics of VBench
@@
In the second approach, we categorize the evaluation metrics based on their correlation, as illustrated in Figure 8 for the image-to-text generation and Figure 9 for the text-to-video generation. In the image-to-video generation, several evaluation metrics are quite specific and possess uniqueness, such as metrics for OCR and coding tasks, leading to a higher number of categories upon classification. In contrast, VBench’s evaluation metrics can easily be divided into several broad classes, which include static video metrics like subject and background consistency, dynamic metrics such as the dynamic degree, video quality indicators like aesthetic and imaging quality, as well as the degree of match between the generated video and the prompt, including object class and spatial relationship. The dynamic degree is particularly noteworthy as it negatively correlates with many other metrics and combats those that prefer static videos, thus preventing videos with no movement from being rated as optimal. Based on these observations, for the text-to-video generation, we apply a hierarchical clustering algorithm to divide VBench metrics into three categories based on their correlations, roughly classified into three types: static video metrics, dynamic video metrics, and video quality along with video-text matching. Subsequently, we select the best-performing OP for each of these three metric categories, where each excels in different aspects. These selected OPs are then combined and experimented with together.
B.3 Setting and Ablation Study Based on T2V-Turbo
@@
The results shown in Table 2 represent the enhancements we achieved based on T2V-Turbo (Li et al., 2024). T2V-Turbo applies LoRA (Hu et al., 2021) to VideoCrafter-2.0 (Chen et al., 2024b) and is trained on the WebVid (Bain et al., 2021) dataset, using VideoCrafter-2.0 as a teacher for distillation and incorporating reinforcement learning with rewards for the generated videos. We make the following cumulative modifications to T2V-Turbo:
@@
  • Proposed Data Pool. Utilizing the optimal data pool identified in Section 5.3, which consists of 147k instances with low video NSFW scores and high frame-text similarities, we replace the WebVid data and train for 6 epochs based on the insight from Section 5.3.
@@
  • Initialization for LoRA. We use the T2V-Turbo model to initialize the parameters for training with LoRA.
@@
  • Self-Distillation. When we initialize with the T2V-Turbo model, the student model is already outperforming the teacher model, VideoCrafter-2.0, potentially leading to unstable training. To mitigate this, we use the T2V-Turbo model itself as the teacher to ensure that the reinforcement learning does not diverge excessively.
@@
  • Real-Data Loss. To enhance the role of the proposed data pool during training, we add a real-data loss between the generated videos and the input videos to the distillation loss and reward loss. Furthermore, we set the weights of both the real-data loss and the distillation loss to 0.5.
@@
Table 5 presents the ablation experiments of our modifications in the VBench evaluation. It is clear that when we replace the WebVid data with our proposed data pool, the model experiences a notable improvement, with the total score increasing from 81.01% to 81.84%. Subsequently, initializing the LoRA training parameters with the T2V-Turbo model does not lead to further enhancements in model performance. We suspect this might be because the teacher model is less effective than the T2V-Turbo model. Therefore, we use the T2V-Turbo model for self-distillation. While this method effectively raises the semantic score, it results in unstable video generation characterized by significant temporal flickering, which severely lowers the video quality. To counteract this, we add a real-data loss with the input data to secure the quality of the generated videos. Ultimately, while improving the semantic score, we ensure the video generation quality remains high and establish a new state-of-the-art.

### v2

@@ D.2 Evaluation Metrics
where $N$ is the number of involved metrics, $s_{i}$ is the score of $i$-th model measurement metric, $s^{\prime}_{i}$ is the corresponding score gained by the baseline model trained on randomly sampled data. Below are the specific evaluation metrics involved in this study.
@@
Across all experiments, results are reported as averages with standard deviations from 2 to 5 repetitions for the image-to-text and text-to-video tasks, respectively, due to their differing levels of variance.
Metrics for Image-Text Pre-training task. For the CLIP experiments, we adopt 40 distinct benchmark scores used in DataComp (Gadre et al., 2023), e.g., the image classification on 30 data subsets for diverse scenarios like ImageNet derivatives, the image-text retrieval on 3 data subsets like Flicker and MSCOCO, the fairness related classification task like Dollar Street and GeoDE. For more details about these tasks, please refer to the DataComp and their original papers.
D.3 Descriptions of Studied Operators
@@
image_text _similarity_filter text, image CLIP Image-Text Similarity Image-text feature cosine similarity based on a CLIP model
@@
The study involves 31 OPs from Data-Juicer (Chen et al., 2024a). Their corresponding statistics and detailed descriptions are provided in Table 4.
D.4 Image-to-Text Generation
Our first task focuses on foundational image understanding ability, by experimenting on Mini-Gemini (MGM-2B), a state-of-the-art (SOTA) 2 billion parameter multimodal LLM (Li et al., 2024b). The training protocol for MGM-2B involves two stages: pretraining and fine-tuning. Our experimental focus lies in the pretraining phase, which seeks to harmonize visual and textual representations. We utilize the original pretraining dataset as our original dataset $\mathcal{D}$, consisting of approximately 1.2M instances. We set the size of $\mathcal{D}_{sample}$ as 200k. The single-OP data pools $\mathcal{D}_{i}$ and multi-OP data pools $\mathcal{D}_{\mathcal{S}}$ are capped at a maximum of 200k instances, ensuring consistency of data pool size. To match the down-sampling rate used during pretraining, the fine-tuning dataset is sampled into a 240k instance subset.
We first conduct single-OP experiments (Section 4.2) that encompasses 22 text-image relevant OPs from Data-Juicer, split evenly between text-only and image-related multimodal OPs. After the two-stage training, model evaluation is conducted on established benchmarks including TextVQA (Singh et al., 2019), MMBench (Liu et al., 2023), and MME (Fu et al., 2023).
For multi-OP data pools (Section 4.3), we identify the top-3 highest-performing OPs from single-OP experiments and study their possible combinations. Additionally, we analyze the correlations among the 23 data statistics produced by 22 OPs capable of generating instance-level stats  [^2] The image height and width are produced by one OP.. Employing a hierarchical clustering algorithm  (Ward Jr, 1963), these OPs are grouped into three clusters based on correlation coefficients, with the highest-performing OP from each cluster selected for combination testing. To ensure a robust and fair comparison, we must acknowledge the constraints imposed by the limited data volume within the highest-tier data pool. As the number of combinations increases, the available dataset size diminishes. In particular, the size of $\mathcal{P}_{\mathcal{S}}$ was reduced from 200k samples to 159k samples during the Top-3 combination experiments. Similarly, in the cluster-wise combination experiments, the dataset size decreased from 200k samples to 126k samples.
Next, in Section 4.4, we explore the optimal OP combination based on previous experiments and adopt the methodology from Section 3.2.3 for comparative experiments on training with repeated data versus non-repeated data. Note that due to filtering, the final instance count decreases from 200k to approximately 159k after the OP combination. These data are then repeated in increments from double to eightfold, mirroring the size of the original pretraining set.
Collectively, all these experiments yield profound insights into image-to-text model training, data processing, and iteration strategies from a data-model co-development perspective, further verified in the larger-scale scenario in Section 4.5.
In terms of the model training details, we train the MGM-2B model from scratch with less training data (about 1/6 of the original training datasets) in baseline experiments to make sure each experiment can be finished within one day. We keep every training setting (e.g. learning rate scheduler, global batch size) the same as the original model except for training datasets and training devices. For single-OP and OP combination experiments are trained on only 1 A100 GPU for each experiment so we increase the number of gradient accumulation steps from 4 to 32 to keep the same global batch size. For experiments of duplicating high-quality datasets, 8 A100 GPUs are involved to train the model, and the number of gradient accumulation steps is restored to 4. Each experiment is repeated 3 times with different random seeds to make the final results more comprehensive.
D.5 Text-to-Video Generation
For the second task, text-to-video generation, we adopt the advanced DiT-based models, EasyAnimate (Xu et al., 2024b), which originally integrates diverse datasets totaling 1.2M instances from InternVid (Wang et al., 2023b) (606k), Panda-70M (Chen et al., 2024e) (605k), and MSR-VTT (Xu et al., 2016) (6k). The studied baseline model is trained on a subset of 40k instances, employing LoRA (Hu et al., 2021) for efficiency. As a result, the size of $\mathcal{D}$ is 1.2M, and the size of $\mathcal{D}_{sample}$, the single-OP data pools $\mathcal{D}_{i}$ and multi-OP data pools $\mathcal{D}_{\mathcal{S}}$ are all 40k. Model outputs are assessed using VBench (Huang et al., 2024) across 16 metrics on video quality and video-text matchness.
Our investigation covered 21 OPs, including 13 text-only OPs and 10 video-related multimodal OPs. Analogous to the image-to-text generation, we conduct single-OP and multi-OP combination experiments, in Section 4.2 and Section 4.3 respectively. However, given the reduced relevance of data statistics in video-related OPs, our analysis centers on the correlations among the 16 VBench evaluation metrics. These metrics are clustered into three groups, with the best-performing OP selected from each group.
Through OP combination experiments, we pinpoint the most effective set of OPs. In Section 4.4, we then sample 40k instances from the filtered data pool and repeat the training process for up to 10 epochs. For comparative analysis, we adhere to the method outlined in Section 3.2.3, selecting larger data volumes (80k, 120k, …, up to 400k instances) for single-epoch training. To examine the effectiveness of the derived insights for text-to-video data-model co-development, we finally incorporate them into larger-scale scenarios in Section 4.5.
In terms of the model training details, we adopt the advanced DiT-based EasyAnimate (Xu et al., 2024b) model, which integrates diverse datasets totaling 1.2M instances from InternVid (Wang et al., 2023b) (606k), Panda-70M (Chen et al., 2024e) (605k), and MSR-VTT (Xu et al., 2016) (6k). Baseline experiments are executed on a subset of 40k instances, employing LoRA (Hu et al., 2021) for efficiency. During training, we maintain a video resolution of 256x256, sample every other frame, and randomly select sequences of 16 consecutive frames. The training process involves performing a backward pass for the loss of every 8 samples, with single-OP and OP combination experiments trained on a single GPU with a batch size of 8 for 5k steps, amounting to approximately 16 GPU hours per training run. Experiments for duplicating high-quality data, as well as larger-scale training, are conducted with a batch size of 1 across 8 GPUs. The models employ the Adam optimizer for training, with a learning rate set to $2\times 10^{-5}$, weight decay parameter at $3\times 10^{-2}$, and epsilon configured to $10^{-10}$. Each experiment is repeated twice with random seeds of 42 and 45, respectively.
D.6 Image-Text Pre-training
For the third task, image-text pre-training, we adopt the well-studied CLIP model (Radford et al., 2021). Specifically, we utilize data from the small track of the DataComp competition (Gadre et al., 2023) and adhere to its evaluation metrics, which include 40 distinct evaluation subsets. Due to some broken links, we successfully downloaded 85.2% of the dataset, resulting in a total of 10.9 million samples as our $\mathcal{D}$. All baseline models were trained on an equivalent volume of data as used in the contrastive experiments, sampled randomly from this dataset.
In the multi-OP experiments (see Fig. LABEL:fig:recipe_clip_top-3), due to data limitations in the top-level data pool, the training dataset size is reduced to 880k for the top recipes. Correspondingly, the default computation scale of DataComp is adjusted to 0.25 expansion rate for the subsequent compute scaling experiments (see Fig. LABEL:fig:scaling-lines).
Appendix E Additional Experimental Results
E.1 Summarization of Sandbox Experiments varying Multiple Scales
Table 5: Overview of sandbox scenarios and their effective evidence across different scales.
{tblr}
width = colspec = Q[140]Q[260]Q[280]Q[269], hlines, vlines, Sandbox Scenario & Image-to-Text Generation Text-to-Video Generation Image-Text Pretraining
Main Effectiveness Evidence Optimal recipe derived from small data pools achieves superior model performance in larger data pools, using only half of baseline’s compute cost. Insights from small data pools results in VBench-Top1 model, transferring across data size, model scales, and model architectures Best recipe identified in the model with fewer FLOPS maintains optimal performance with increased model FLOPS and compute resources, showing clear scaling behaviors in power-law form.
Model Scale Range MGM-2B in all experiments EasyAnimate to T2V-Turbo
(Heterogeneous architectures) CLIP: ViT-B-32 to ViT-B-16
(Different FLOPS)
Data Scale Range
(w.r.t Distinct Dataset Size)
126k to 200k 40k to 147k and 228k 880k to 2,683k
Compute Scale Range (w.r.t Number of Trained Sample) 1 to 8 Epochs 1 to 10 Epochs 4 to 14 Epochs
The table 5 provides an overview of our sandbox experiments on image-to-text generation, text-to-video generation, and image-text pre-training tasks.
E.2 Complete Single-Operator Ranking
@@
Table 7: The complete OP ranking, including their statistical dimensions and the improvements relative to the baseline for image-text similarity/classification task. We consider three splits with low, middle, and high statistical values for each OP. The baseline used is based on random sampling with equal data volume.
OP-Generated Statistics Average Performance Changes (%)
Data Pool (Low) Data Pool (Mid) Data Pool (High)
CLIP Image-Text Similarity -32.57 -6.39 39.53
BLIP Image-Text Similarity -24.28 1.82 25.39
Image NSFW Score 12.18 1.28 -18.38
Word Number -18.65 0.74 9.78
Stopword Ratio -4.28 -3.33 8.97
Special Character Ratio 8.86 4.15 -16.03
Phrase Grounding Recall 7.79 1.85 -10.60
Text Length -8.31 1.81 7.29
Character Repetition Ratio 1.99 0.04 6.63
Image Aspect Ratio 4.93 -4.55 5.87
Text Perplexity 5.27 2.46 -9.56
Image Width -6.66 4.97 5.23
Image Height -4.03 5.02 0.89
Image Size -12.11 5.00 2.87
Image Aesthetics Score -9.61 -8.13 4.64
Image Watermark Score 3.84 -3.74 -4.72
Flagged Word Ratio 3.66 3.47 1.59
Entity Dependency Number -5.53 -0.39 2.50
Word Repetition Ratio -3.16 -0.91 1.84
Alphanumeric Ratio -2.55 1.65 0.63
Token Number -6.35 1.44 0.27
In Table 6 and Table 7, we present complete numeric results conducted on individual OP experiments (Section 4.2), from which we can discern some more detailed observations.
In image-to-text generation, it is preferable for the input of training images to align as closely as possible with the original configuration of the vision tower, such as training dimensions (height, width, and sizes). Additionally, CLIP similarity scores tend to be more influential than BLIP similarity scores. The BLIP similarity does not show much distinction and paradoxically, a lower similarity often results in better performance, which defies common sense. Images with excessively high aesthetic quality may offer limited assistance in feature alignment, while watermarks might have certain impacts on the OCR performance of the model.
@@
In image-text pre-training, we can see that the top three performing operations are image_text_similarity_filter, image_text_matching_filter, and image_nsfw_filter. The first two OPs generate statistics based on auxiliary CLIP and BLIP models, enlightening that the modality alignment degree is critical to the studied pair-wise learning task.
E.3 Diversity Analysis
In this subsection, we delve into the diversity of the data in interested data pools. Here, we confine our focus to statistical analysis of words within text data and compute their entropy. We operate under the assumption that the texts provide an accurate description of the images and videos. Consequently, the diversity inferred from the texts also serves as a proxy for the diversity of the associated images and videos.
Table 8: The entropy of text words for data pools with different levels of Image NSFW score and Language score.
Task OP-Generated Statistics Word Entropy
Data Pool (Low) Data Pool (Mid) Data Pool (High)
Image-to-Text Image NSFW Score 6.97 7.35 7.29
Language Score 7.47 7.32 6.98
Text-to-Video Video NSFW Score 5.84 6.03 6.01
Language Score 6.30 5.85 5.73
The Table 8 shows the entropy of text words for different data pools, which can be normalized as follows:
$\sum_{w}-\mathcal{P}(w)\log{\mathcal{P}(w)},$ (2)
where $w$ is a word and $\mathcal{P}$ is the distribution of words in a data pool. As we can see, data pools with higher NSFW scores and lower language scores have higher word entropy, suggesting greater diversity within these data pools.
E.4 Correlation Analysis
@@
To investigate the intrinsic relationships between OPs and to support our recipe formulation, we explore relevance from the following two perspectives.
@@
Figure 8: Pearson correlation coefficients for OP statistics in image-to-text generation.
@@
Figure 9: Pearson correlation coefficients for OP statistics in text-to-video generation.
@@
First, we adopt the most direct approach by examining the Pearson correlation coefficients between the statistics of OPs, as illustrated in Figures 8 and 9. Intuitively, the associations between the statistics of OPs utilized in image-to-text generation appear to be significantly stronger than those in text-to-video generation. For instance, in image-to-text generation, phrase grounding recall shows a strong positive correlation with text perplexity and special character ratio, while exhibiting a strong negative correlation with the alphanumeric ratio, language score, number of text actions, stopword ratio, and text length. In contrast, in text-to-video generation, we observe relationships primarily among the purely textual OPs, while video-related operators are largely orthogonal to others. Therefore, for image-to-text generation, we categorize OPs into three groups based on the correlation of their statistics and select the optimal OP from each category to create combinations. However, this method does not appear appropriate for text-to-video generation due to the sparser correlations observed.
@@
Figure 10: Pearson correlation coefficients for each dimension in TextVQA, MMBench, and MME.
Figure 11: Pearson correlation coefficients for evaluation metrics in VBench.
In the second approach, we categorize the evaluation metrics based on their correlations, as illustrated in Figure 10 for image-to-text generation and Figure 11 for text-to-video generation. In image-to-text generation, several evaluation metrics are highly specific and possess unique characteristics, such as those for OCR and coding tasks, resulting in a greater number of categories upon classification. In contrast, VBench’s evaluation metrics can be broadly divided into several classes, including static video metrics (e.g., subject and background consistency), dynamic metrics (e.g., dynamic degree), and video quality indicators (e.g., aesthetic quality and imaging quality).
@@
Notably, the dynamic degree negatively correlates with many other metrics, particularly those favoring static videos, thus preventing videos without movement from being rated as optimal. Based on these observations, for text-to-video generation, we apply a hierarchical clustering algorithm (Ward Jr, 1963) to classify the VBench metrics into three categories based on their correlations: static video metrics, dynamic video metrics, and video quality along with video-text matching. We then select the best-performing OP for each of these three metric categories, where each excels in different aspects. These selected OPs are subsequently combined and experimentally evaluated together.
E.5 Recipes Based on Correlation Analysis
Figure 13: The improvements (%) of recipes with different OP combinations. The cluster recipes stem from the combinations of the best OPs in 3 categories. In the image-to-text generation, due to the limited data volume in the top-level data pool, the size of $\mathcal{P}_{\mathcal{S}}$ decreased from 200k to 126k in the cluster recipes.
In addition to selecting the overall best-performing OPs, we aim to identify operators with distinct advantages to explore whether combining these operators can synergistically leverage their strengths for improved outcomes. We selected operators with unique strengths based on correlation information obtained from single-operator experiments. Detailed correlation analyses can be found in Appendix E.4.
Specifically, for the image-to-text generation, we categorize the OPs by calculating correlations between their statistics. We then select representative operators within each category: TextActionFilter for text, ImageNSFWFilter for images, and PhraseGroundingRecallFilter for combined text-image relevance. For the text-to-video generation, based on single-operator experiments, we categorize the metrics of VBench (Huang et al., 2024) into three classes according to their relevance and select the best-performing operator from each category. The chosen operators are VideoMotionScoreFilter, which excels in video static features; VideoDurationFilter, which is superior in dynamic features; and VideoAestheticsFilter, which exhibits the best composite performance in video quality and video-text matching. We summarize the experiment results in Figures LABEL:fig:recipe_image-to-text_cluster-3 and LABEL:fig:recipe_text-to-video_cluster-3.
Observation 8 (Effect of Orthogonal Combination) Combining OPs that excel in orthogonal dimensions on model or data does not guarantee complementary effects; rather, it is more likely that they will impede each other’s performance.
As depicted in Figures LABEL:fig:recipe_image-to-text_cluster-3 and LABEL:fig:recipe_text-to-video_cluster-3, regardless of how these top-performing OPs are combined, they ultimately reduce the model’s performance in both image-to-text and text-to-video generation. This observation challenges the naive assumption widely used in existing SOTA works, that various intuitively useful data cleansing actions, when stacked serially, can synergistically enhance performance.
E.6 Detailed Evaluation Results on MGM
The detailed experimental results in terms of different evaluation benchmarks are shown in Table 9.
Table 9: Performance of MGM-2B with different pretraining datasets. MGM-2B pretrained with only 1/10 distinct instances and 1/2 of the total instances performs better. than the baseline trained on the full dataset. The “*” indicates our reproduced version that is comparable with the official version.
MGM-2B Num. of Instances Avg. Perf. Changes (%) MMBench MMBench-CN MME-Perception MME-Cognition
Baseline* 1226k - 59.2 51.6 1334 302
Data-Juicer 159k (x4) +2.12 62.08 52.32 1323.37 311.07
E.7 Ablation Study Based on T2V-Turbo
@@
The results are shown in Table 3 representing the enhancements we achieved based on T2V-Turbo (Li et al., 2024a). T2V-Turbo applies LoRA (Hu et al., 2021) to VideoCrafter-2.0 (Chen et al., 2024d) and is trained on the WebVid (Bain et al., 2021) dataset, using VideoCrafter-2.0 as a teacher for distillation and incorporating reinforcement learning with rewards for the generated videos. The loss $L$ of T2V-Turbo is defined as:
$L=L_{CD}-1\times R_{img}-2\times R_{vid},$ (3)
where $L_{CD}$ is the loss of consistency distillation (Song et al., 2023), $R_{img}$ is the reward of image quality of generated video frames from HPSv2.1 (Wu et al., 2023) and $R_{vid}$ is the reward of video quality from InternVideo2 (InternVid2 S2) (Wang et al., 2024b). In the paper, we make the following cumulative modifications to T2V-Turbo:
  1. 1.
@@
    Proposed Data Pool (Enhancement from Data View). Utilizing the optimal data pool identified in Section 4.4, which consists of 147k instances with low video NSFW scores and high frame-text similarities, we replace the WebVid data and train for 6 epochs based on the insight from Section 4.4.
  2. 2.
@@
    Initialization for LoRA (Enhancement from Model View). We use the T2V-Turbo model to initialize the parameters for training with LoRA.
  3. 3.
@@
    Self-Distillation. (Enhancement from Model View). When we initialize with the T2V-Turbo model, the student model is already outperforming the teacher model, VideoCrafter-2.0, potentially leading to unstable training. To mitigate this, we use the T2V-Turbo model itself as the teacher to ensure that the reinforcement learning does not diverge excessively.
  4. 4.
@@
    Real-Data Loss (Data-Model Co-Enhancement). To enhance the role of the proposed data pool during training, we add a real-data loss between the generated videos and the input videos to the distillation loss and reward loss. Furthermore, we set the weights of both the real-data loss and the distillation loss to 0.5. The modified loss from Equation 3 can be specified as:
    $L=0.5\times L_{CD}+0.5\times L_{real}-1\times R_{img}-2\times R_{vid},$ (4)
    where ×$\times L_{real}$ is the real-data loss. Training on 147k instances with 6 epochs, we obtain the model, Data-Juicer (T2V, 147k).
  5. 5.
    Self-Evolution (Data-Model Co-Enhancement). To validate the scalability of our methodologies, we expanded the original data pool and, by applying the video_aesthetics_filter and video_motion_score_filter to the existing recipes, we select an additional 228k instances to evolutionary self-distill Data-Juicer (T2V, 147k). Meanwhile, in order to further amplify the impact of data, we have reduced the emphasis on reinforcement learning, thereby:
    $L=0.5\times L_{CD}+0.5\times L_{real}-0.2\times R_{img}-0.4\times R_{vid}.$ (5)
    We finally get Data-Juicer (DJ, 228k) in 4 epochs training.
@@
Table 10 presents the ablation experiments of our modifications in the VBench evaluation. It is clear that when we replace the WebVid data with our proposed data pool, the model experiences a notable improvement, with the total score increasing from 81.01% to 81.84%. Subsequently, initializing the LoRA training parameters with the T2V-Turbo model does not lead to further enhancements in model performance. We suspect this might be because the teacher model is less effective than the T2V-Turbo model. Therefore, we use the T2V-Turbo model for self-distillation. While this method effectively raises the semantic score, it results in unstable video generation characterized by significant temporal flickering, which severely lowers the video quality. To counteract this, we add a real-data loss with the input data to secure the quality of the generated videos. Moreover, we evolve the model by using our trained model as the teacher model for continuous training on additional data, rather than T2V-Turbo. Ultimately, as we continue to enhance both the quality and semantic scores, we establish a new state-of-the-art.
E.8 Full Results on VBench leaderboard
Figure 14: The VBench Leaderboard as of September 23, 2024, illustrating the rank 1 achievement empowered by our data-model co-development workflow.
Table 11: Leading models on the VBench leaderboard as of Sep 23, 2024. “Board Avg.” denotes the weighted average of normalized scores across 16 metrics defined by VBench. “Quality Avg.” represents the aggregated scores from 7 video quality metrics, whereas “Semantic Avg” aggregates scores from 9 metrics evaluating the consistency between prompts and generated videos. “Uniform Avg.” indicates the simple average of scores related to quality and semantic metrics. Data-Juicer (T2V, 147k) is based on T2V-Turbo distillation training on 147k instances. In order to demonstrate the scalability of our methodologies, we then further self-distilled our model to obtain Data-Juicer (DJ, 228k) on other 228k instances.
1. Data-Juicer (DJ, 228k) 82.53 81.26 83.38 79.13
   Data-Juicer (T2V, 147k) 82.10 80.54 83.14 77.93
2. Gen-3 (RunwayML, 2024) 82.32 79.64 84.11 75.17
3. VEnhancer (VC2) (He et al., 2024a) 81.97 80.00 83.27 76.73
4. Kling (2024-07) (Kuaishou, 2024) 81.85 79.54 83.39 75.68
5. LaVie-2 (Wang et al., 2023a) 81.75 79.50 83.24 75.76
6. CogVideoX-5B (Yang et al., 2024) 81.61 79.90 82.75 77.04
7. Vchitect 2.0-2B (Vchitect, 2024) 81.57 80.15 82.51 77.79
8. T2V-Turbo (VC2) (Li et al., 2024a) 81.01 78.67 82.57 74.76
Table 12: Detailed FLOPs comparison for representative T2V models. The factor $\alpha$ indicates a
Models Performance Full size of used Dataset Inference FLOPS per Sample (Estimated) Number of Training Samples (Estimated) Training Cost in EFLOPS
Ours
(stems from VC2) 82.53 228k 12T/sample, 320p * 8fps 640k $7.68\alpha$
Ours
(stems from VC2) 82.10 147k 12T/sample, 320p * 8fps 640k $7.68\alpha$
VEnhancer
(stems from VC2) 81.97 350k 478T/sample, 720p * 24fps $\geq$ 350k $\geq 167.3\alpha$
CogVideoX-5B 81.61 1,051M 338T/sample, 720p * 8/16fps 35M $\geq 11841\alpha$
Vchitect 2.0-2B 81.57 - 252T/sample, 768p * 8fps - -
Backed by our proposed methodology and experiments, our Data-Juicer (DJ, with 228k) model refreshes the state-of-the-art on VBench (Huang et al., 2024), surpassing models like Gen-3 (RunwayML, 2024) and Kling (Kuaishou, 2024), as illustrated in Figure 14 and detailed in Table 3.
E.9 Cost Comparison of all Sandbox Experiments
Besides, we present the compute details for our experiments and compare them to other baselines. Our experiment on a single data pool is designed to be completed within one day using a GPU. Consequently, with 8 A100 GPUs, we can run all our experiments within approximately 3 days, 7 days, and 10 days for the CLIP, image-to-text, and text-to-video experiments, respectively. The training sample size is in the order of $O(1)$ million video-text pairs and $O(1)$ million image-text pairs in our experiments. As a reference, the pre-training cost for a cutting-edge text-to-video model like Meta’s MovieGen involves 6,144 H100 GPUs, $O(100)$ million video-text pairs, and $O(1)$ billion image-text pairs.
More specifically, for the image-to-text experiments, in Table 2, our MGM model shows a +2.12% average performance improvement over baseline, using approximately 51.8% of the FLOPS.
For the CLIP experiments, refer to Figure 12; the FLOPs with full expansion rate are 189 $\alpha$ PFLOPS (14.78G inference FLOPS per sample) and 526 $\alpha$ PFLOPS (41.09G inference FLOPS per sample) for B-32 and B-16, respectively.
For the text-to-video experiments, Noting that estimating the FLOPS for some models in Table 3 and Table 8 poses several challenges due to the lack of publicly available information. For example, some of their training codes, and training datasets have not been open-sourced, and specific details such as the resolution and frame rate of training videos are unavailable. Thus, we report the estimated FLOPS for the models with sufficient details about the training samples and open-sourced models below. We utilized the inference code of their released open-source models and measured the FLOPS during inference using the calflops  [^3] https://github.com/MrYxJ/calculate-flops.pytorch tool. We then estimated the training FLOPS by applying a factor of $\alpha$ and taking into account the disclosed quantity of training data. From this we can see that our method achieves strong performance while utilizing fewer FLOPS, especially in comparison to the third-place model, VEnhancer, which also originates from VC2, the same upstream model as ours. The second-place Gen-3 and the fourth-place Kling are omitted as they are proprietary models within commercial products.